OpenStudy (anonymous):

If A and B are independent events, P(A) = 0.45, and P(B) = 0.25, find the probabilities below P(AuB)

jimthompson5910 (jim_thompson5910):

P( A u B ) = P(A) + P(B) - P(A intersect B) P( A u B ) = 0.45 + 0.25 - 0.45*0.25 P( A u B ) = 0.5875

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

You're welcome.